OpenStudy (nicholausblackmon):

hey! i have a question! do you know how to find the midpoint of P(-2,8) and Q(8,-4)? Please help me. Could you please walk me through the formula?! I would love to give you a medal if help is given!

OpenStudy (nicholausblackmon):

@dan815

OpenStudy (nicholausblackmon):

(3,2) (3,3) (2,2) (2,3) these are the answers

OpenStudy (anonymous):

\[x_{midd} = \frac{ (x_1 + x_2) }{ 2 }\] \[y_{midd} = \frac{ (y_1 + y_2) }{ 2 }\]

OpenStudy (nicholausblackmon):

so i just add the point they are at so if x was at -4 i would put it as -4-n?

OpenStudy (nicholausblackmon):

just as writing the equation.

OpenStudy (anonymous):

\[x_{midd} = \frac{ -2+8 }{ 2 } = ?\]

rvc (rvc):

add coordinates of x term

OpenStudy (nicholausblackmon):

oh ok

rvc (rvc):

as told by @sangya21

OpenStudy (nicholausblackmon):

=3?

OpenStudy (anonymous):

Correct. So what will be the y?

rvc (rvc):

correct :)

OpenStudy (nicholausblackmon):

give me a second please

OpenStudy (nicholausblackmon):

6?

OpenStudy (nicholausblackmon):

how would it be 6?

OpenStudy (nicholausblackmon):

@dan815

OpenStudy (anonymous):

\[y_{midd} = \frac{ 8+(-4) }{ 2 } = \frac{ 8-4 }{ 2 }\]

OpenStudy (nicholausblackmon):

look at the second y

OpenStudy (nicholausblackmon):

you would be subracting a negative so it would technically be adding the 4 to it would be 12?

OpenStudy (nicholausblackmon):

OHHH ADD

OpenStudy (dan815):